### v1

# Fréchet Power-Scenario Distance: A Metric for Evaluating Generative AI Models across Multiple Time-Scales in Smart Grids

Abstract
Generative artificial intelligence (AI) models in smart grids have advanced significantly in recent years due to their ability to generate large amounts of synthetic data, which would otherwise be difficult to obtain in the real world due to confidentiality constraints. A key challenge in utilizing such synthetic data is how to assess the data quality produced from such generative models. Traditional Euclidean distance-based metrics only reflect pair-wise relations between two individual samples, and could fail in evaluating quality differences between groups of synthetic datasets. In this work, we propose a novel metric based on the Fréchet Distance (FD) estimated between two datasets in a learned feature space. The proposed method evaluates the quality of generation from a distributional perspective. Empirical results demonstrate the superiority of the proposed metric across timescales and models, enhancing the reliability of data-driven decision-making in smart grid operations.
Index Terms:
Generative model, synthetic data evaluation, machine learning, energy data
I Introduction
Generative models in the electric energy sector have been an active field of research in the past few years, thanks to their potential to create realistic and diverse scenarios for system planning, reliability assessment, and renewable energy integration—ultimately enhancing grid resilience and operational efficiency. These models, such as Generative Adversarial Networks (GANs), allow researchers to access much larger sets of synthetic data across multiple time scales that would otherwise be unavailable due to confidentiality constraints [1]. In contrast to traditional methods that involve creating synthetic power networks and subsequently using commercial-grade simulation software to generate electrical measurement variables [2], these generative approaches leverage a data-driven methodology. They produce large volumes of synthetic electrical data by utilizing historical electrical measurements in a computationally efficient manner.
While these generative models hold great potential for a wide range of power system applications, a central question remains largely unexplored: What would be a fair and robust metric to evaluate the quality of various generative techniques in creating synthetic power system data across multiple time resolutions? Conventionally, evaluation methods are based on Euclidean metrics between pairs of samples (signals), such as Mean Squared Error (MSE) and Mean Absolute Percentage Error (MAPE) [3, 4, 5, 6]. However, these methods are susceptible to the selection of samples, leading to inconsistent evaluations. Moreover, the sample-by-sample comparison approach poses challenges when evaluating datasets produced by generative models, where the distributional properties of the datasets are more critical.
As an alternative, distribution-based comparison methods, such as Fréchet Distance (FD), which can capture the broader distributional characteristics of generative models, have been explored in previous work [7]. However, applying FD directly at the data level may not effectively capture the similarity between datasets. Other approaches have opted for task-specific metrics that involve statistical analysis of task-based parameters. For instance, the number of zeros is employed as a task-specific indicator to assess the quality of voltage sag generation in [8]. Similarly, the indicators used to detect anomalies in electricity consumption include residential stimulus allocation, residential tariff allocation, and SME (Small and Medium-sized Enterprise) allocation in [9]. However, these approaches introduce inconsistencies, as the evaluation criteria vary across tasks. The variability makes it challenging to achieve a unified understanding of a model’s overall performance or to establish standardized benchmarks, potentially leading to biased assessments.
Building on insights from computer vision research, where generative model evaluation has been extensively explored, we investigate methods that could establish more reliable evaluation standards. One such approach is the Inception Score (IS) [10], which was initially adopted to assess the quality and diversity of generated images using a pre-trained Inception network. However, IS faced criticism for its inability to detect the mode collapse of the generative model, where the model produces a limited variety of outputs or concentrates on only a few modes within the distribution, resulting in potential failures across various scenarios [11]. The Fréchet Inception Distance (FID) was proposed to overcome these limitations. FID computes the Fréchet distance between the distributions of real and generated images in a vectorized feature space mapped by a pre-trained Inception network, namely Inception v3. This metric considers both the mean and covariance of the feature space of real and generated data distributions, making it more sensitive to variations in both quality and diversity while closely aligning with human perception [12]. [13] suggests using Kernel Inception Distance (KID), which implements Maximum Mean Discrepancy (MMD) with kernel functions applied to the feature space of the Inception model can further improve the metric performance compared to FID. Unlike FID, which can suffer from estimator bias. KID provides an unbiased estimate, ensuring stable evaluation, particularly for small sample datasets. However, this approach is highly dependent on the choice of kernel functions, with different kernels potentially leading to inconsistent results.
Inspired by the success of the FID metric, in this paper we aim to develop a robust metric that could work in the context of multi-time-scale generative data sets in smart grids. The challenge that arises in adapting FID for this context are as follows: (1) The original FID feature extractor, Inception v3, is optimized primarily to capture image-specific features, making it potentially unsuitable for power system data; (2) FID metric is developed based on the ImageNet dataset, which contains over 14 million images across 1,000 categories as a comprehensive benchmark for most image generation tasks [14], whereas there is no standardized dataset of comparable scope and diversity for power systems; (3) Inception V3 model requires a fixed input resolution of 299x299 pixels, which is incompatible with the multi-resolution and multi-duration characteristics inherent to power system data.
To address these three challenges, we (1) train a new feature extraction model to capture both spatial and temporal features critical for power system data, enhancing its performance on such datasets; (2) collect comprehensive data set which, to the best of our knowledge, captures the major synthetic tasks associated with power systems; and (3) design a hierarchical structure composed of multiple feature extractors to handle multi-resolution datasets, enabling features from shorter timescales to be effectively integrated into models operating at longer timescales.
The key contributions of this paper are suggested as follows:
  1. 1)
    A hierarchy multi-resolution feature extraction network: We present a novel network design where multiple feature extractors are trained to handle data at different time resolutions and are connected in a hierarchical structure. This approach enables data of varying resolutions to be accommodated within the model, ensuring that features from shorter timescales are effectively transferred to longer timescales, thereby enhancing the model’s ability to capture the complex temporal dynamics inherent in power systems.
  2. 2)
    A task-independent metric for generative models in smart grid: To our best knowledge, this is first-of-its-kind introduction of the Fréchet Power-Scenario Distance (FPD) in the smart grids context. FPD is specifically developed and trained to assess the performance of generative models in the power system domain. It can effectively handle multi-timescale data and provides a unified standard for a diverse range of power system tasks.
The rest of the paper is organized as follows: Section II details the design of the proposed evaluation metrics; Section A introduces model training detail; Section IV demonstrates the effectiveness of FPD across various disturbances and its advantages over traditional metrics; and Section V concludes the paper.
II Design of Evaluation Metrics
II-A Overview of the feature extraction model
We employ neural networks to extract high-level features capable of capturing complex spatial and temporal relations. Currently, a general feature extractor for this task does not exist in the power system domain. Implementing such a general model is challenging due to the diverse time resolutions and durations inherent in power system data. To address this challenge, we build a unified feature extraction framework that accommodates such diversity. As shown in Fig.1, our proposed framework is a multi-resolution multi-duration model, where a series of feature extractors, each trained for a specific time resolution, are interconnected hierarchically. This design allows multiple extractors to be combined for different data durations, with output features from each higher-resolution extractor feeding into the input of the subsequent lower-resolution extractor. This integration of high-resolution features into lower-resolution modules allows for effective utilization of detailed data in lower-resolution models without significantly increasing dimensional complexity.
II-B Hierarchical feature extraction model
Considering the different timescales underlying the steady-state and transient-state data in power systems, we proposed two types of feature extraction models respectively. Due to the different sampling intervals, there are multiple resolutions and durations in steady-state data. Therefore, we adopt a multi-layer hierarchical architecture. For the transient-state situations, the sample intervals are much simpler, typically a selection of integer multiples of the rated frequency. Hence, we designed a single-layer model for the transient-state case.
The steady-state model can process time-series data with different resolutions $\{r_{s}\}$. The resolution progresses sequentially as follows:
[…]
To accomodate steady-state data in different resolutions and durations, we employ a multi-layer structure. Each layer is designed to process data at a specific resolution $r_{s}$ and transform it into a feature space aligned with the next higher resolution $r_{s+1}$. We denote the extractor handling data at resolution $r_{s}$ as $\mathcal{M}_{r_{s}}(x_{r_{s}})$, where $x_{r_{s}}$ represents data from the $r_{s}$ resolution. Each encoding function $\mathcal{M}_{r_{s}}$ takes an input $x_{r_{s}}$ The extractor then outputs a feature representation $z_{r_{s}}$ for the next resolution $r_{s+1}$ model processing
We employ these extractors in different resolutions in a hierarchical structure, where each subsequent extractor builds upon the features aggregated from the stack of extractors in the prior level. A switch mechanism is introduced so that each extractor can adapt to the type of input it receives. Each extractor can select its input from either the features extracted at the prior resolution level, or the raw data at the current resolution level, as demonstrated in Fig.1. For instance, the daily module ($\mathcal{M}_{r_{1}}$) can directly process 24 hourly data points if only hourly data is available, producing hourly features on a daily scale. Alternatively, if only 5-minute data is available, 24 hourly modules ($\mathcal{M}_{r_{0}}$) can be used to transform the 5-minute data for each hour (consisting of 12 data points per hour) into corresponding hourly features. Subsequently, the resulting 24 hourly features are concatenated and utilized as input to the daily module.
It is noteworthy that this structure has the capability to adapt multi-resolution, multi-duration data. The input data can be injected at any designated entry point in the model, depending on the data resolution $r_{s}$. The model will process the data through a series of lower-level extractors to higher-level extractors and finally generate the desired feature representation at the module corresponding to the target scale $r_{e}$.
During training and usage, the model processes the data in the form $\left(\frac{N}{L_{r_{s}}},D_{r_{s}},L_{r_{s}}\right)$ given the resolution of data $r_{s}$. Here, $N$ denotes the total number of data points across all sequences. For example, consider 1,000 data samples, each representing daily-scale data recorded at 5-minute intervals. Since one day has $288$ such intervals, the total number of data points is $1{,}000\times 288=288{,}000$ $L_{r_{s}}$ is the number of data points required at resolution $r_{s}$ to form a single data point at the next higher resolution $r_{s+1}$; for example, for 5-min resolution data, $L_{\text{5-min}}=12$. Therefore, to process the entire batch of sequences, the dataset must be divided into segments of size $\frac{N}{L_{r_{s}}}$. Under this setup, $\frac{N}{L_{r_{s}}}$ modules $\mathcal{M}_{r_{s}}$ will each handle a segment of length $L_{r_{s}}$, which aligns with the module design in our structure. Next, we define $D_{r_{s}}$ as the augmented feature dimension associated with resolution $r_{s}$. This dimension includes features extracted from the preceding module, as well as a one-dimensional mean-value derived from the prior-level data. In contrast, the raw time-series data remain one-dimensional. In this case, the higher-level data is viewed as the mean-value of the lower-level data values, and the other lower-level extracted features can be simply zero-padded. This dimension, $D_{r_{s}}$, needs to be chosen properly so that feature space distribution is approximately Gaussian distributed because the subsequent Fréchet Distance calculation assumes that the features follow a Gaussian distribution.
Mathematically, to extract feature vectors for input data $x_{r_{s}}$ at resolution $r_{s}$ over a target duration $r_{e}$, we define the feature extraction process as a sequential operation within the hierarchical framework. The feature output at each resolution can be expressed as:
$z_{r_{s}}=\mathcal{M}_{r_{s}}\left(\mathbf{I}_{r_{s}}\right),$ (1)
where $\mathbf{I}_{r_{s}}$ indicates the input at level $r_{s}$:
$\mathbf{I}_{r_{s}}=\begin{cases}\left[z_{r_{s-1}},\bar{x}_{r_{s-1}}\right],&\text{if $z_{r_{s-1}}$ is available}\\ \left[0,x_{r_{s}}\right],&\text{if $x_{r_{s}}$ is available}\\ \end{cases}$
In each resolution, we first determine input $\mathbf{I}_{r_{s}}$ between the feature representation from the prior level and the data at the current level. Due to the dimensional inconsistency between $\left[z_{r_{s-1}},\bar{x}_{r_{s-1}}\right]$, which has dimension $D_{r_{s}}$, and $x_{r_{s}}$, which has dimension $1$, we need to pad $x_{r_{s}}$, if necessary, to match the dimension $D_{r_{s}}$. This operation ensures that all choices of $\mathbf{I}_{r_{s}}$ have the same dimension and can be adapted by the model. Also, considering the days in different months are inconsistent, for $r=2$, we pad the length to 31 if needed. Then we reshape the $\mathbf{I}_{r_{s}}$ to size $\left(\frac{N}{L_{r_{s}}},D_{r_{s}},L_{r_{s}}\right)$. This operation allows each module at a given resolution to process only a segment of the entire input sequence. Then, by stacking multiple such modules together, we ensure that the entire sequence in $r_{e}$ scale is effectively processed. The reshaped feature vector $z_{r_{s}}$ is passed as input to the extractor at the resolution $r_{s+1}$. We then update the resolution to $r_{s}=r_{s+1}$ and repeat this process sequentially until target resolution $r_{e}$ is reached. Once is $r_{e}$ achieved, we output the resulting feature vectors, which can be used to calculate the FPD score.
The parameters and modules corresponding to each resolution $r$ are listed in Table I. The feature extraction algorithm is outlined in Algorithm 1.
TABLE I: Parameters and modules in steady-state model
$i$ $r_{i}$ $D_{r_{i}}$ $L_{r_{i}}$ $\mathcal{M}_{r_{i}}$
0 5-min 1 12 Hourly Module
1 Hourly 10 24 Daily Module
2 Daily 14 31 Monthly Module
3 Monthly 20 12 Yearly Module
4 Yearly
Algorithm 1 Feature Extraction Algorithm
1:  Input: Duration $r_{e}$, Start resolution $r_{s}$, Batch size $N$, Input data $x_{r_{s}}$, Feature dimension $D_{r_{s}}$, Data Length $L_{r_{s}}$ at resolution $r_{s}$
2:  Output: Final feature representation for data $x_{r_{s}}$ in $r_{e}$ scale
3:  while $r_{s}<r_{e}$ do
4:     if $r_{s-1}$ is available then
5:        Set $\mathbf{I}_{r_{s}}$ to $\left[\mathbf{z}_{r_{s-1}},\bar{x}_{r_{s-1}}\right]$
6:     else if $r_{s}$ is available then
7:        Set $\mathbf{I}_{r_{s}}$ to $[0,x_{r_{s}}]$
8:     end if
9:     Reshape ${I}_{r_{s}}$ to size $\left(\frac{N}{L_{r_{s}}},D_{r_{s}},L_{r_{s}}\right)$
10:     Calculate feature representation: $z_{r_{s}}=\mathcal{M}_{r_{s}}\left(\mathbf{I}_{r_{s}}\right)$
11:     Advance resolution: $r_{s}=r_{s+1}$
12:  end while
13:  Return Output final feature vector $z_{r_{s}}$
The architecture designs are simpler in the transient-state cases. For this case, we use a single-time resolution with a 120Hz sampling frequency. Hence, we only apply a single feature extractor to the transient-state related task. A data sample is an 8-second length sequence with three channels of voltage magnitude, phase angle, and frequency, all represented in per-unit form. The input is first upsampled to 32 channels and finally downsampled by stride of 2 filters to a feature vector with 2048 dimensions. Then, we can calculate the FPD scores based on the extracted feature vectors.
For the architecture design of a single feature extractor in both models, we utilize residual blocks in the feature extractor as demonstrated in Fig.1. The residual connections have been widely studied and proved to increase neural network depth to enhance representation capabilities. We utilize the conv-1d and batchnorm-1d layers for accommodating the time-series inputs. This Resnet-like feature extractor is a base component for different scenarios and timescale tasks, followed by simple linear heads for different downstream training objectives.
Figure 1: Overview of the hierarchical multi-resolution feature extraction framework, integrating multiple time-scale feature extractors to enhance generative model evaluation in power systems.
II-C Fréchet Power-Scenario Distance (FPD) calculation
Once robust representations are obtained from the feature extractor, we can then compare the distribution similarity between the two datasets in the feature space. Three distribution evaluation metrics have often been adopted[7]: (1) KL-divergence, (2) Kernel Distance (Maximum Mean Discrepancy - MMD), and (3) Fréchet Distance (Wasserstein-2 Distance).
The features can often be approximately viewed as having a Gaussian distribution. Denote the feature vectors extracted from two sets $x_{r}^{1}$ and $x_{r}^{2}$ at resolution $r$ as $z_{r}^{1}$ and $z_{r}^{2}$. The means of extracted features can then be expressed as:
$\mathbf{m}_{1}=\frac{1}{N}\sum_{i=1}^{N}z_{r,i}^{1},\quad\mathbf{m}_{2}=\frac{1}{N}\sum_{i=1}^{N}z_{r,i}^{2}$ (2)
The covariance matrices of the feature vectors extracted are:
$\Sigma_{1}=\frac{1}{N}\sum_{i=1}^{N}\left(z_{r,i}^{1}-\mathbf{m}_{1}\right)\left(z_{r,i}^{1}-\mathbf{m}_{1}\right)^{T}$ (3)
$\Sigma_{2}=\frac{1}{N}\sum_{i=1}^{N}\left(z_{r,i}^{2}-\mathbf{m}_{2}\right)\left(z_{r,i}^{2}-\mathbf{m}_{2}\right)^{T}$ (4)
Using the KL-divergence as the evaluating metric, the similarity between $z_{r}^{1}$ and $z_{r}^{2}$ can be assessed as [15]:
$\text{KL}(z_{r}^{1}||z_{r}^{2})=$ $\frac{1}{2}\bigg{(}\text{Tr}(\Sigma_{2}^{-1}\Sigma_{1})+(\mathbf{m}_{2}-\mathbf{m}_{1})^{T}\Sigma_{2}^{-1}(\mathbf{m}_{2}-\mathbf{m}_{1})$ (5)
$-d+\log\left(\frac{\det(\Sigma_{2})}{\det(\Sigma_{1})}\right)\bigg{)},$
where $d$ is the dimensionality of the feature space $z_{r}^{1}$ and $z_{r}^{2}$. It is important to note that when applying the KL-divergence to compare the distance between two normal distributions, both $z_{r}^{1}$ and $z_{r}^{2}$ must have the same dimensionality. Because our goal is to design a metric that can compare datasets in different resolutions which have different feature dimensions, the KL-divergence is not suitable for our purpose. It is also well known that the KL divergence has the disadvantage of being asymmetric and unstable in some cases.
With the Kernel Distance as the evaluation metric, the distance between $z_{r}^{1}$ and $z_{r}^{2}$ is calculated as:
$\text{MMD}(z_{r}^{1},z_{r}^{2})=$ (6)
$\;\|\mathbf{m}_{1}-\mathbf{m}_{2}\|^{2}+\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\Big{(}k(z_{r,i}^{1},z_{r,j}^{1})$
$\;+k(z_{r,i}^{2},z_{r,j}^{2})-2k(z_{r,i}^{1},z_{r,j}^{2})\Big{)},$
where $k(x,y)$ is a kernel function (e.g., Gaussian kernel). Choosing a proper kernel is extremely important for the metric to be effective, as different kernel functions can lead to varying metric performance.
We propose to use the Fréchet Distance due to its stability and consistency as a metric based on our experiments. Since this metric is specifically tailored to evaluate power system scenarios, we name it Fréchet Power-Scenario Distance (FPD). The $\text{FPD}(z_{r}^{1},z_{r}^{2})$ between the extracted features at resolution $r$ is calculated as:
$\text{FPD}(z_{r}^{1},z_{r}^{2})=\|\mathbf{m}_{1}-\mathbf{m}_{2}\|^{2}+\text{Tr}(\Sigma_{1}+\Sigma_{2}-2(\Sigma_{1}\Sigma_{2})^{1/2})$ (7)
FPD is essentially a form of the Fréchet Distance, where we assume that the features of the two datasets follow Gaussian distributions. This metric quantifies the similarity between the two distributions: a smaller FPD value indicates a more significant similarity between the feature distributions, ultimately reflecting a higher similarity between the datasets.
III Model Training
The training data used in this study consists of multiple open-source datasets, each capturing different generative tasks. These datasets include various types of demand, EV charging profiles, and various types of electricity generation. Each dataset varies in terms of resolution and timescale, contributing to a comprehensive representation of the power system domain. The specific characteristics of each dataset are summarized in Table II.
TABLE II: Summary of Datasets Used for Training
Dataset Description Source Resolution Timescale
1 Wind Generation[16] NREL 5-minute 1 year
2 Solar Generation[17] NREL 5-minute 1 year
3 EV charging profiles[18] Northwest Power and Conservation Council Hourly 1 year
4 Commercial user load[19] OpenEI.org Hourly 1 years
5 Residential user load[19] OpenEI.org Hourly 1 years
6 Load Data in Texas[20] ERCOT Hourly 3 years
7 Demand by balancing authority[21] EIA daily 3 years
8 Generation by balancing authority[21] EIA daily 3 years
9 Utility Scale Electricity Net Generation[22] EIA monthly 14 years
10 Phasor Measurement Unit PSSE simulations 0.016-second(60Hz) 50 hours
For steady-state model, each module in the feature extraction model is trained individually and sequentially from the bottom level (5-min module) to the top level (yearly module). Once a certain level module has finished training, the parameters will be fixed, and we start to train the next level module. The training data for the module at each resolution is $I_{r_{s}}$ The training dataset is shown in Table II
Each module is trained based on the loss function:
$L_{r_{s}}=\frac{1}{N}\sum_{i=1}^{N}\Bigg{[}$ $\sum_{k=1}^{K}\left(f_{\text{reg}}^{k}(I_{r_{s}})-y_{i}^{\text{reg}}\right)^{2}$ (8)
$-\sum\limits_{i=1}^{C}y_{i}^{\text{clf}}\log\left(\text{softmax}(f_{\text{clf}}(I_{r_{s}}))\right)\Bigg{]}$
Both $f_{\text{reg}}$ and $f_{\text{clf}}$ are fully connected layers that take the feature output from the module and generate the corresponding predictions of tasks. Here, $W$ denotes the layer weights, and $b$ denotes the layer biases:
$f_{\text{reg}}^{k}(I_{r_{s}})=W_{\text{reg}}^{k}\mathcal{M}(I_{r_{s}})+b_{\text{reg}}^{k}$ (9)
$f_{\text{clf}}(I_{r_{s}})=W_{\text{clf}}\mathcal{M}(I_{r_{s}})+b_{\text{clf}}$ (10)
The first term in the (8), $\sum_{k=1}^{K}\left(f_{\text{reg}}^{k}(I_{r_{s}})-y_{i}^{\text{reg}}\right)^{2}$, corresponds to the Mean Squared Error (MSE) loss across a total of $K=9$ regression tasks. Each regression task is designed for leading extractors to capture features that reflect a specific statistical characteristic of the data, represented by $y_{i}^{\text{reg}}$. The list of regression tasks is shown in Table V in Appendix. The second term,$-\sum\limits_{i=1}^{C}y_{i}^{\text{clf}}\log\left(\text{softmax}(f_{\text{clf}}(I_{r_{s}}))\right)$, represents the Cross-Entropy Loss for the classification task which categorize data based on classification labels $y_{i}^{\text{clf}}$. $y_{i}^{\text{clf}}$ is determined by the dataset index shown in Table II and $C$ is the number of categories of dataset. The hyper-parameters for training are shown in Table VI in Appendix.
For transient-state model training, we employ the fault type, amplitude minimum, and maximum as supervised pretraining labels. We then exploit the pre-trained feature extractor to evaluate synthetic samples. The length of a data sequence is 960, and input data is encoded into a 2048-dimension representation. Moreover, the hyper-parameters for training are shown in Table VII in Appendix. Once training is complete, the tasks are irrelevant, and the obtained module $\mathcal{M}(I_{r_{s}})$ will be used for feature extraction.
All training and experiments were conducted on a 16 GB Apple M2 chip MacBook Pro. The model was trained using PyTorch 2.2, leveraging Apple’s Metal Performance Shaders (MPS) backend for GPU acceleration.
IV Effectiveness Evaluation of the FPD
IV-A Effectiveness on different disturbances
(a) Gaussian Noise
(b) Missing Data
(c) Solar Data Contaminate
(d) Gaussian Smooth
(e) Error Cumulate
(f) Time Shift
Figure 2: Effect of six different types of disturbances on the dataset $X$: (2(a)) Gaussian Noise is added with mean 0 and varying variances $\alpha=[0,0.16,1.6,4]$, where a larger $\alpha$ indicates more noise. (2(b)) Missing Data is simulated by randomly replacing $\alpha\times 100\%$ of the points in each sample of $X$ with zeros, where $\alpha=[0,0.1,0.25,0.5]$. (2(c)) Solar Data Contamination replaces $\alpha\times 100\%$ of the points in $X$ with corresponding points from a solar power dataset $Y$[17], where $\alpha=[0,0.25,0.5,0.75]$. (2(d)) Gaussian Smooth applies Gaussian filter which has variance $\alpha=[0,10,20,30]$ to the original dataset. Where a larger $\alpha$ indicates more significant smoothing. (2(e)) Error Cumulate is simulated by adding an error, which is times with a random Gaussian noise with a mean of 1 and varying variances $\alpha=[0,0.005,0.01,0.03]$ at each time-step. (2(f)) Time Shift is a potential problem that exists in time-series generation. To simulate this effect, we shift the data forward by $\alpha=[0,40,60,80]$ intervals.
In this section, we conduct several experiments to demonstrate the effectiveness of FPD. The experiments are designed based on the Wind Generation dataset 1 in Table II. We progressively increase the disturbance applied to the original dataset X in each experiment to provide a controllable and scalable measure of FPD performance.
  • Gaussian Noise: In this case, Gaussian noise with a mean of 0 and a variance $\alpha$ is added to $X$. Where $\alpha\in[0,0.16,1.6,4]$ represents the disturbance added. Larger $\alpha$ indicates more noise added.
  • Missing Data: To simulate missing data points in $X$, we randomly replace some values with zeros for each sample in this case. The disturbance $\alpha\in[0,0.1,0.25,0.5]$ indicates that $\alpha\times 100\%$ of the points in $X$ are randomly set to zero.
  • Solar Data Contamination: To show FPD can distinguish the data from another dataset, we contaminate the wind dataset $X$ with data from the solar power dataset $Y$ [17]. The disturbance $\alpha\in[0,0.25,0.5,0.75]$ indicates that $\alpha\times 100\%$ of the samples in $X$ are replaced by corresponding points from $Y$.
  • Gaussian Smooth: To simulate scenarios where the generative model may fail to capture fine details of the original data, resulting in overly smooth outputs, we intentionally apply a Gaussian filter to blur these details. This operation allows us to test whether the FPD metric can detect such changes. The disturbance parameter $\alpha\in[0,10,20,30]$ represents the standard deviation of the Gaussian filter applied to the original data, with higher values of $\alpha$ indicating more robust smoothing.
  • Error Accumulation: Certain generative models, like Recurrent Neural Networks (RNNs), struggle with long sequences due to cumulative error over time. We introduce an accumulating error by applying Gaussian noise at each time step to simulate this effect. Specifically, for each time step $t$ in sequence length $T$, we add a noise factor $\epsilon_{t}^{(\alpha)}\sim\mathcal{N}(1,\alpha)$ which is sampled from a Gaussian distribution centered at 1 with varying standard deviations, $\alpha=[0,0.005,0.01,0.03]$. For each time step, we compute cumulative modification factors $E^{(\alpha)}_{t}$ recursively:
    $E^{(\alpha)}_{t}=E^{(\alpha)}_{t-1}\cdot\epsilon_{t}^{(\alpha)},\quad E^{(\alpha)}_{0}=1$
    The modified error cumulated data $\tilde{x}^{(\alpha)}$ is then given by:
    $\tilde{x}^{(\alpha)}_{t}=x_{t}\cdot E^{(\alpha)}_{t}$
    where $x_{t}$ is the original data at time $t$. We demonstrate that the FPD metric effectively detects and highlights this error accumulation effect.
  • Time Shift: Time shift in time-series data is a common issue in generative models, often leading to misalignment of temporal features between the original and generated data. This misalignment can be particularly problematic in power system data, where certain temporal features, such as peaks, carry critical information. To simulate this, we shift the data forward by intervals of $\alpha=[0,40,60,80]$, creating misleading temporal dependencies. Our results demonstrate that FPD effectively detects and captures these temporal shifts.
As shown in Fig. 2, KL-divergence can be inconsistent due to its asymmetry feature, while Kernel Distance may fail to capture significant changes and is sensitive to kernel choice. Additionally, both methods also exhibit scale variability across disturbances. In contrast, Fréchet Distance consistently reflects disturbance levels and maintains stable scaling, making it a more reliable evaluation metric. For these reasons, we select Fréchet Distance as the evaluation metric and refer to our method as Fréchet Power-Scenario Distance (FPD).
Figure 3: Sample-wise distance MAPE VS. FPD
IV-B Comparison of FPD with conventional evaluation metrics
Figure 4: Data-level Fréchet distance (FD) VS. FPD
IV-B1 Sample-Wise Euclidean Distance (MAPE) vs. Distribution-Wise Distance (FPD)
Traditional Euclidean-based metrics such as Mean Absolute Percentage Error (MAPE) typically perform sample-wise comparisons, making the results highly dependent on the selection of sample pairs between the two datasets. This approach can lead to inconsistent and potentially invalid outcomes. To illustrate this, we revisit the Gaussian noise case presented in Fig. 2(a). In this example, we make MAPE to match sample pairs randomly between the original and noise-perturbed datasets. As shown in Fig. 3, in some instances, higher noise levels result in lower MAPE values, which contradicts the ground truth. In contrast, the proposed FPD metric demonstrates expected behavior, providing a more effective measurement.
IV-B2 Data-Level Distribution Comparison (Fréchet Distance) vs. Feature-Level Distribution Comparison (FPD)
As previously noted, the data-level Fréchet Distance (FD) metric is often ineffective at capturing high-level complex patterns. FPD metric addresses this issue by incorporating feature extraction and measuring the similarity between distributions in the feature space. Feature extraction enables a better representation of higher-level patterns and structures within the data, allowing FPD to capture more relevant similarities for evaluating the quality of generated data.
To demonstrate this advantage, we constructed a fabricated dataset by randomly sampling from a distribution with a similar mean and covariance to the original data. Despite these similarities, the fabricated dataset lacks the temporal patterns present in the original dataset, resulting in a sparse relation. As shown in Fig. 4, while FD fails to identify this disparity, FPD accurately reflects the weaker relationship between the fabricated and original datasets.
IV-C Coherence between FPD and the downstream tasks
To assess whether FPD’s indication of high-quality data can be used to determine the practical usefulness for downstream tasks, we further evaluate its efficacy for transient-state case using PMU data containing three types of faults. We employ the Brownian noises to mimic the accumulating temporary errors in the PMU data prediction as depicted in Fig. 5. As shown in Table III, we cross-evaluate different fault scenarios. Hence, we can observe non-zero diagonal elements in the FPD score matrix, along with the diagonal values being significantly lower than the non-diagonal elements, showing the capability of identifying different scenarios other than evaluating generation quality. To further prove the capability of FPD can handle the downstream task in steady-state case, we reuse the wind dataset[16] and define the sub-task as data classification based on the average power across the day as used in Chen’s work [23]. The three cases differ based on average power levels, as shown in Fig. 6. Moreover, the cross-evaluation of different mean value scenarios is shown in the Table.IV again highlights the alignment between the FPD evaluation and the specific sub-task.
Figure 5: Ground-truth and synthetic PMU voltage magnitude curves under different fault scenarios. Synthetic data is modulated using Brownian noises.
TABLE III: Cross-evaluation on FPD performance for PMU data.
Synthetic $|V|$ Ground-truth $|V|$ bus_trip branch_trip branch_fault
bus_trip 744.26 1086.07 2586.15
branch_trip 1462.24 846.03 2720.04
branch_fault 2516.35 2401.13 687.99
Figure 6: Ground-truth and synthetic wind power under different mean scenarios. Synthetic data is modulated using Brownian noises.
TABLE IV: Cross-evaluation on FPD performance for wind power data.
Synthetic $|MW|$ Ground-truth $|MW|$ $\bar{x}<0.5$ $\bar{x}<3$ $\bar{x}>6$
$\bar{x}<0.5$ 736.97 1300.42 901.36
$\bar{x}<3$ 1300.42 565.35 2720.04
$\bar{x}>6$ 901.36 989.83 686.51
IV-D Cross-model comparison with FPD
In the previous sections, all evaluations were conducted directly on the artificially generated datasets. This approach enables precise control over generation quality and model performance assessment. However, FPD can also be used to evaluate model generated performing the same task, as demonstrated in the following two scenarios.
  1. 1.
    Scenario 1: EV Charging Profile Data Generation
    Accurately evaluating the quality of synthetic data in real-world settings presents a significant challenge, as highlighted by [24], for which FPD provides a solution. [24] proposed a novel generative model, DiffCharge, based on a diffusion model designed to conditionally generate electric vehicle (EV) charging power profiles. Probability density function (PDF) comparisons and discriminative scores derived from a trained classifier were used to evaluate battery-level data generation performance between DiffCharge and baseline models such as Time-GAN. However, evaluations for station-level data generation were not given. We conduct experiments to compare DiffCharge’s performance with a W-GAN model on the station level; see Fig. 7. We find that DiffCharge achieves a lower FPD score. Visually, we also observe that DiffCharge can capture the general trend better, while W-GAN generates less realistic peaky signals. Therefore, our results indeed support the conclusion that the DiffCharge model can more accurately approximate the distribution of EV charging profiles at both the station-level and battery-level.
  2. 2.
    Scenario 2: Wind Power Data Generation
    To further demonstrate FPD’s cross-model comparison ability, we replicate the study from [1], where the authors introduced a model called StyleGAN, which integrates forecast and climate data to enhance the generation of renewable energy data. They showed that StyleGAN is superior to WGAN proposed in [23] based on multiple uncertainty quantification and statistical metrics, such as RMSE. In our experiment, the differences between generations from StyleGAN and WGAN were challenging to distinguish through visual inspection. StyleGAN’s results seemed to capture the overall trend more accurately. As shown in Fig. 8, FPD confirms StyleGAN aligns better with real data, supporting the findings in [1].
    Figure 7: Scenario 1 – Comparison of DiffCharge and W-GAN in generating EV charging profiles at the station level.
    Figure 8: Scenario 2 – Comparison of W-GAN and StyleGAN in generating wind power data.
IV-E Comparability across varying time resolutions
A key feature of FPD is its ability to adapt to varying time resolutions within a single metric. This flexibility enables effective comparisons across datasets with different time resolutions. To demonstrate this capability, we collected hourly resolution data for solar and wind power and evaluated their similarity with the 5-minute resolution solar and wind power datasets used in Section IV-A. The results, shown in Fig. 9, indicate that FPD successfully identifies that the solar power datasets are more similar to each other than they are to the wind power data despite their differing time resolutions. This result demonstrates the metric’s effectiveness in facilitating comparability across datasets with different temporal resolutions.
Figure 9: The FPD metric is evaluated between the left:5-min resolution solar power dataset[17], right top:hourly resolution solar power dataset[25] and right bottom:hourly resolution wind power dataset[25]. The results demonstrate that FPD effectively generalizes across different time resolutions.
V Conclusion and Future work
This paper introduces the Fréchet Power-Scenario Distance (FPD) as a novel, task-agnostic metric for evaluating generative model outputs in the power systems sector. FPD addresses the limitations of traditional metrics by employing a hierarchical feature extraction framework that captures multi-resolution temporal information. Then, the Fréchet Distance is applied to assess the quality of generated data in the feature space. This method enables FPD to provide an accurate, robust, and consistent evaluation across diverse power system applications.
Comprehensive case studies demonstrate FPD’s effectiveness in distinguishing data quality across various disturbances, time resolutions, and sub-task scenarios. These results underscore FPD’s potential as a standardized cross-model approach for validating generative models, offering support for both enhancing model quality and promoting wider adoption within the power systems domain.
There are some potential studies for future work. 1) Currently, the dataset we are using is still limited due to accessibility constraints. Thus, building a more comprehensive dataset tailored to a wider range of tasks within the power system domain is possible to improve the performance of FPD; 2) While the current model structure can handle multi-resolution and multi-duration data, it is not yet sufficiently generalized to accommodate data of any resolution or duration. Therefore, there is potential for developing a more generalized structure that can address this challenge and further enhance the versatility of the metric.
Appendix A Appendix: Feature extraction training detail
TABLE V: Regression labels for steady-state training
Regression Tasks
Minimum
Maximum
Range
Variance
Number of data points above average
Number of data points below average
Skewness
Kurtosis
Coefficient of variation
TABLE VI: Hyper-parameters for steady-state model training
Parameter Setting
Minibatch size 1024
Optimizer Adam
Optimizer: learning rate 1e-3
TABLE VII: Hyper-parameters for transient-state model training
Parameter Setting
Minibatch size 128
Optimizer Adam
Optimizer: learning rate 1e-4
Optimizer: weight decay 2e-5
Classification loss coefficient 1.0
Regression loss coefficient 1.0

### v2

@@
-Fréchet Power-Scenario Distance: A Metric for Evaluating Generative AI Models across Multiple Time-Scales in Smart Grids
+Fréchet Power-Scenario Distance: A Metric for Evaluating Synthetic Data Across Multiple Time-Scales in Smart Grids
 Abstract
@@
 A key challenge in utilizing such synthetic data is how to assess the data quality produced from such generative models.
-Traditional Euclidean distance-based metrics only reflect pair-wise relations between two individual samples, and could fail in evaluating quality differences between groups of synthetic datasets.
+Traditional metrics such as sample-wise Euclidean distance and distributional distances applied directly to raw generated data inadequately reflect higher-order temporal dependencies and cross-temporal relationships between real and synthetic series, and thus struggle to discriminate generative quality.
 In this work, we propose a novel metric based on the Fréchet Distance (FD) estimated between two datasets in a learned feature space.
-The proposed method evaluates the quality of generation from a distributional perspective.
-Empirical results demonstrate the superiority of the proposed metric across timescales and models, enhancing the reliability of data-driven decision-making in smart grid operations.
-Index Terms:
-Generative model, synthetic data evaluation, machine learning, energy data
+The proposed method assesses synthetic data quality via distributional comparisons in a feature space derived from a model tailored to the smart grid domain.
+Empirical results demonstrate the superiority of the proposed metric across downstream tasks and generative models, enhancing the reliability of data-driven decision-making in smart grid operations.
 I Introduction
@@
 Moreover, the sample-by-sample comparison approach poses challenges when evaluating datasets produced by generative models, where the distributional properties of the datasets are more critical.
-As an alternative, distribution-based comparison methods, such as Fréchet Distance (FD), which can capture the broader distributional characteristics of generative models, have been explored in previous work [7].
-However, applying FD directly at the data level may not effectively capture the similarity between datasets.
+As an alternative, distribution-based approaches—covering proper scoring rules such as the Continuous Ranked Probability Score (CRPS) [7] and Energy Score (ES)[8], and distribution-to-distribution distances such as Kullback–Leibler (KL) divergence [9], Maximum Mean Discrepancy (MMD)[10], and Wasserstein distance [11]—have been explored in prior work to capture the broader distributional characteristics of synthetic datasets.
+However, when computed directly on raw data, distribution-level metrics tend to underweight higher-order structures such as temporal dependencies, which are crucial in smart grid time-series.
+Consequently, they can fail to discriminate meaningful changes in generative quality that primarily manifest in the dependence structure [12].
 Other approaches have opted for task-specific metrics that involve statistical analysis of task-based parameters.
@@
 However, these approaches introduce inconsistencies, as the evaluation criteria vary across tasks.
-The variability makes it challenging to achieve a unified understanding of a model’s overall performance or to establish standardized benchmarks, potentially leading to biased assessments.
+Such variability makes it difficult to form a unified view of overall synthetic data quality or to establish standardized benchmarks, potentially leading to biased assessments.
 Building on insights from computer vision research, where generative model evaluation has been extensively explored, we investigate methods that could establish more reliable evaluation standards.
-One such approach is the Inception Score (IS) [10], which was initially adopted to assess the quality and diversity of generated images using a pre-trained Inception network.
+One such approach is the Inception Score (IS) [15], which was initially adopted to assess the quality and diversity of generated images using a pretrained Inception network.
 However, IS faced criticism for its inability to detect the mode collapse of the generative model, where the model produces a limited variety of outputs or concentrates on only a few modes within the distribution, resulting in potential failures across various scenarios [16].
 The Fréchet Inception Distance (FID) was proposed to overcome these limitations.
-FID computes the Fréchet distance between the distributions of real and generated images in a vectorized feature space mapped by a pre-trained Inception network, namely Inception v3.
-This metric considers both the mean and covariance of the feature space of real and generated data distributions, making it more sensitive to variations in both quality and diversity while closely aligning with human perception [12].
-[13] suggests using Kernel Inception Distance (KID), which implements Maximum Mean Discrepancy (MMD) with kernel functions applied to the feature space of the Inception model can further improve the metric performance compared to FID.
-Unlike FID, which can suffer from estimator bias.
-KID provides an unbiased estimate, ensuring stable evaluation, particularly for small sample datasets.
-However, this approach is highly dependent on the choice of kernel functions, with different kernels potentially leading to inconsistent results.
-Inspired by the success of the FID metric, in this paper we aim to develop a robust metric that could work in the context of multi-time-scale generative data sets in smart grids.
-The challenge that arises in adapting FID for this context are as follows: (1) The original FID feature extractor, Inception v3, is optimized primarily to capture image-specific features, making it potentially unsuitable for power system data; (2) FID metric is developed based on the ImageNet dataset, which contains over 14 million images across 1,000 categories as a comprehensive benchmark for most image generation tasks [14], whereas there is no standardized dataset of comparable scope and diversity for power systems; (3) Inception V3 model requires a fixed input resolution of 299x299 pixels, which is incompatible with the multi-resolution and multi-duration characteristics inherent to power system data.
-To address these three challenges, we (1) train a new feature extraction model to capture both spatial and temporal features critical for power system data, enhancing its performance on such datasets; (2) collect comprehensive data set which, to the best of our knowledge, captures the major synthetic tasks associated with power systems; and (3) design a hierarchical structure composed of multiple feature extractors to handle multi-resolution datasets, enabling features from shorter timescales to be effectively integrated into models operating at longer timescales.
-The key contributions of this paper are suggested as follows:
+FID evaluates distributional similarity by taking the Fréchet (squared 2-Wasserstein) distance between the Gaussian embeddings of real and generated images in feature spaces extracted by a pretrained Inception network.
+[17]
+Inspired by the success of the FID metric, we argue that computing a distributional metric (e.g., Fréchet distance) in a learned feature space yields a more comprehensive and fair evaluation of smart grid synthetic datasets.
+The central challenge is thus obtaining a valid feature extractor that provides features aligned with the smart grid characteristics.
+Inception-v3, the encoder used by FID, is optimized for images rather than time-series.
+Thus, when applied to smart grid data, it underrepresents temporal dynamics.
+Moreover, its fixed $299\times 299$ input requirement is incompatible with the inherently multi-resolution, multi-duration nature of power-system signals.
+A natural alternative is to consider time-series encoders explicitly designed to capture temporal dynamics and to be flexible with respect to input length.
+Although prior work has produced generic time-series encoders (e.g., TS2Vec [18], TimeMixer [19]), they are not plug-and-play for smart grids.
+While these encoders can handle variable sequence lengths—typically by padding or truncating to a fixed window—their learned representations often emphasize broad statistical or sequential regularities and miss physics-driven structure.
+In contrast to generic time-series, smart grid signals exhibit distinctive structure: strong diurnal and seasonal periodicities, hard physical constraints (e.g., ramp limits, nonnegativity, no solar at night), and a multi-resolution character.
+Capturing these attributes is crucial for downstream tasks such as anomaly detection, disturbance classification, and evaluation of generative scenarios, yet generic pretrained encoders frequently fail to preserve them.
+To address this gap, we propose a power-scenario-aware feature extractor that explicitly encodes the temporal and spectral signatures of grid data.
+Leveraging a multi-resolution hierarchical architecture and trained on large-scale power datasets, the extractor uses a classification head to emphasize local dynamics/pattern regimes and a regression head to capture broader statistical trends.
+Together, these objectives yield representations that faithfully capture the physical underpinnings of power-system data.
+Computing a distributional distance (e.g., Fréchet) in this feature space then provides a more insightful, effective, and robust evaluation of power-domain generative models.
+The key contributions of this paper are as follows:
   1. 1)
-    A hierarchy multi-resolution feature extraction network: We present a novel network design where multiple feature extractors are trained to handle data at different time resolutions and are connected in a hierarchical structure.
-This approach enables data of varying resolutions to be accommodated within the model, ensuring that features from shorter timescales are effectively transferred to longer timescales, thereby enhancing the model’s ability to capture the complex temporal dynamics inherent in power systems.
+    A hierarchical multi-resolution feature extraction network: We present a novel network design where multiple feature extractors are trained to handle data at different time resolutions and are connected in a hierarchical structure.
+This framework accommodates multi-resolution inputs, propagating features learned at shorter timescales to longer horizons.
+Each extractor is trained on smart grid datasets using joint classification and regression tasks, encouraging representations that capture source-specific dynamics as well as general statistical structure.
+Collectively, this improves the model’s capacity to represent the complex temporal behavior inherent in smart grids.
   2. 2)
-    A task-independent metric for generative models in smart grid: To our best knowledge, this is first-of-its-kind introduction of the Fréchet Power-Scenario Distance (FPD) in the smart grids context.
-FPD is specifically developed and trained to assess the performance of generative models in the power system domain.
-It can effectively handle multi-timescale data and provides a unified standard for a diverse range of power system tasks.
-The rest of the paper is organized as follows: Section II details the design of the proposed evaluation metrics; Section A introduces model training detail; Section IV demonstrates the effectiveness of FPD across various disturbances and its advantages over traditional metrics; and Section V concludes the paper.
+    A feature-based, temporally aware distributional metric for generative models in smart grids: Unlike traditional metrics that operate directly on raw data, we propose a metric that evaluates in a learned feature space: specifically, we compute the Fréchet distance between extracted feature embeddings of real and generated data to quantify generative quality.
+Extensive experiments show that this approach robustly and consistently handles multi-timescale data and provides a unified standard across diverse smart grid generative tasks.
+We refer to this metric as the Fréchet Power-Scenario Distance (FPD).
+The rest of the paper is organized as follows: Section II details the design of the proposed evaluation metrics; Section III introduces model training details; Section IV demonstrates the effectiveness of FPD across various disturbances and its advantages over traditional metrics; and Section V concludes the paper.
 II Design of Evaluation Metrics
@@
 We employ neural networks to extract high-level features capable of capturing complex spatial and temporal relations.
-Currently, a general feature extractor for this task does not exist in the power system domain.
-Implementing such a general model is challenging due to the diverse time resolutions and durations inherent in power system data.
+Currently, a general feature extractor for this task does not exist in the smart grid domain.
+Implementing such a general model is challenging due to the diverse time resolutions and durations inherent in smart grid data.
 To address this challenge, we build a unified feature extraction framework that accommodates such diversity.
-As shown in Fig.1, our proposed framework is a multi-resolution multi-duration model, where a series of feature extractors, each trained for a specific time resolution, are interconnected hierarchically.
+As shown in Fig.
+1, our proposed framework is a multi-resolution, multi-duration model, where a series of feature extractors, each trained for a specific time resolution, are interconnected hierarchically.
+Every feature extractor during training is paired with a classification head that emphasizes local dynamics and a regression head that captures broader statistical trends, together ensuring that the extracted features faithfully represent the characteristics of smart grid data.
 This design allows multiple extractors to be combined for different data durations, with output features from each higher-resolution extractor feeding into the input of the subsequent lower-resolution extractor.
-This integration of high-resolution features into lower-resolution modules allows for effective utilization of detailed data in lower-resolution models without significantly increasing dimensional complexity.
+This integration of high-resolution features into lower-resolution modules enables effective use of fine-grained information while avoiding substantial increases in dimensional complexity.
+At the same time, it ensures that the output features at each level remain closely aligned with the characteristic patterns of smart grid datasets.
 II-B Hierarchical feature extraction model
-Considering the different timescales underlying the steady-state and transient-state data in power systems, we proposed two types of feature extraction models respectively.
+Considering the different timescales underlying the steady-state and transient-state data in smart grids, we propose two types of feature extraction models respectively.
 Due to the different sampling intervals, there are multiple resolutions and durations in steady-state data.
 Therefore, we adopt a multi-layer hierarchical architecture.
-For the transient-state situations, the sample intervals are much simpler, typically a selection of integer multiples of the rated frequency.
-Hence, we designed a single-layer model for the transient-state case.
+For the transient-state situations, the sampling intervals are much simpler, typically a selection of integer multiples of the rated frequency.
+Hence, we design a single-layer model for the transient-state case.
 The steady-state model can process time-series data with different resolutions $\{r_{s}\}$. The resolution progresses sequentially as follows:
-[…]
-To accomodate steady-state data in different resolutions and durations, we employ a multi-layer structure.
-Each layer is designed to process data at a specific resolution $r_{s}$ and transform it into a feature space aligned with the next higher resolution $r_{s+1}$.
+$r_{s}\in\{\text{5-min}\rightarrow\text{hourly}\rightarrow\text{daily}\rightarrow\text{monthly}\rightarrow\text{yearly}\}.$
+To accommodate steady-state data in different resolutions and durations, we employ a multi-layer structure.
+Each layer is designed to process data at a specific resolution $r_{s}$ and transform it into a feature space aligned with the following higher resolution $r_{s+1}$.
 We denote the extractor handling data at resolution $r_{s}$ as $\mathcal{M}_{r_{s}}(x_{r_{s}})$, where $x_{r_{s}}$ represents data from the $r_{s}$ resolution.
-Each encoding function $\mathcal{M}_{r_{s}}$ takes an input $x_{r_{s}}$ The extractor then outputs a feature representation $z_{r_{s}}$ for the next resolution $r_{s+1}$ model processing
-We employ these extractors in different resolutions in a hierarchical structure, where each subsequent extractor builds upon the features aggregated from the stack of extractors in the prior level.
+Each encoding function $\mathcal{M}_{r_{s}}$ takes an input $x_{r_{s}}$.
+The extractor then outputs a feature representation $z_{r_{s}}$ for the following resolution $r_{s+1}$ model processing.
+We employ these extractors at different resolutions in a hierarchical structure, where each subsequent extractor builds upon the features aggregated from the stack of extractors in the prior level.
 A switch mechanism is introduced so that each extractor can adapt to the type of input it receives.
-Each extractor can select its input from either the features extracted at the prior resolution level, or the raw data at the current resolution level, as demonstrated in Fig.1.
+Each extractor can select its input from either the features extracted at the prior resolution level, or the raw data at the current resolution level, as demonstrated in Fig.
+1.
 For instance, the daily module ($\mathcal{M}_{r_{1}}$) can directly process 24 hourly data points if only hourly data is available, producing hourly features on a daily scale.
@@
 Subsequently, the resulting 24 hourly features are concatenated and utilized as input to the daily module.
-It is noteworthy that this structure has the capability to adapt multi-resolution, multi-duration data.
+It is noteworthy that this structure has the capability to adapt to multi-resolution, multi-duration data.
 The input data can be injected at any designated entry point in the model, depending on the data resolution $r_{s}$.
@@
 Here, $N$ denotes the total number of data points across all sequences.
-For example, consider 1,000 data samples, each representing daily-scale data recorded at 5-minute intervals.
-Since one day has $288$ such intervals, the total number of data points is $1{,}000\times 288=288{,}000$ $L_{r_{s}}$ is the number of data points required at resolution $r_{s}$ to form a single data point at the next higher resolution $r_{s+1}$; for example, for 5-min resolution data, $L_{\text{5-min}}=12$.
+Consider 1,000 data samples, each representing daily-scale data recorded at 5-minute intervals.
+Since one day has $288$ such intervals, the total number of data points is $1{,}000\times 288=288{,}000.$ $L_{r_{s}}$ is the number of data points required at resolution $r_{s}$ to form a single data point at the next higher resolution $r_{s+1}$.
+For 5-min resolution data, $L_{\text{5-min}}=12$.
 Therefore, to process the entire batch of sequences, the dataset must be divided into segments of size $\frac{N}{L_{r_{s}}}$.
@@
 Next, we define $D_{r_{s}}$ as the augmented feature dimension associated with resolution $r_{s}$.
-This dimension includes features extracted from the preceding module, as well as a one-dimensional mean-value derived from the prior-level data.
+This dimension includes features extracted from the preceding module, as well as a one-dimensional mean value derived from the prior-level data.
 In contrast, the raw time-series data remain one-dimensional.
-In this case, the higher-level data is viewed as the mean-value of the lower-level data values, and the other lower-level extracted features can be simply zero-padded.
-This dimension, $D_{r_{s}}$, needs to be chosen properly so that feature space distribution is approximately Gaussian distributed because the subsequent Fréchet Distance calculation assumes that the features follow a Gaussian distribution.
+In this case, the higher-level data are viewed as the mean value of the lower-level data values, and the other lower-level extracted features can be simply zero-padded.
+The dimension $D_{r_{s}}$ must be selected carefully to ensure that the FPD score computed in this feature space reliably and smoothly reflects the actual differences between datasets with different disturbance levels.
+To determine an appropriate dimension, we perform ablation tests by varying the feature dimension that achieves the best performance across different levels of Gaussian noise disturbance.
+An example ablation result for the hourly module is presented in SI Sec.S1; Fig.
+S1.
+The selected dimensions $D_{r_{s}}$ for all modules are summarized in SI Sec.S1; Table S4.
 Mathematically, to extract feature vectors for input data $x_{r_{s}}$ at resolution $r_{s}$ over a target duration $r_{e}$, we define the feature extraction process as a sequential operation within the hierarchical framework.
@@
 where $\mathbf{I}_{r_{s}}$ indicates the input at level $r_{s}$:
-$\mathbf{I}_{r_{s}}=\begin{cases}\left[z_{r_{s-1}},\bar{x}_{r_{s-1}}\right],&\text{if $z_{r_{s-1}}$ is available}\\ \left[0,x_{r_{s}}\right],&\text{if $x_{r_{s}}$ is available}\\ \end{cases}$
+[…]
 In each resolution, we first determine input $\mathbf{I}_{r_{s}}$ between the feature representation from the prior level and the data at the current level.
@@
 Also, considering the days in different months are inconsistent, for $r=2$, we pad the length to 31 if needed.
-Then we reshape the $\mathbf{I}_{r_{s}}$ to size $\left(\frac{N}{L_{r_{s}}},D_{r_{s}},L_{r_{s}}\right)$.
+We then reshape $\mathbf{I}_{r_{s}}$ to size $\left(\frac{N}{L_{r_{s}}},D_{r_{s}},L_{r_{s}}\right)$.
 This operation allows each module at a given resolution to process only a segment of the entire input sequence.
@@
 The reshaped feature vector $z_{r_{s}}$ is passed as input to the extractor at the resolution $r_{s+1}$.
-We then update the resolution to $r_{s}=r_{s+1}$ and repeat this process sequentially until target resolution $r_{e}$ is reached.
-Once is $r_{e}$ achieved, we output the resulting feature vectors, which can be used to calculate the FPD score.
-The parameters and modules corresponding to each resolution $r$ are listed in Table I. The feature extraction algorithm is outlined in Algorithm 1.
-TABLE I: Parameters and modules in steady-state model
-$i$ $r_{i}$ $D_{r_{i}}$ $L_{r_{i}}$ $\mathcal{M}_{r_{i}}$
-0 5-min 1 12 Hourly Module
-1 Hourly 10 24 Daily Module
-2 Daily 14 31 Monthly Module
-3 Monthly 20 12 Yearly Module
-4 Yearly
+We then update the resolution to $r_{s}=r_{s+1}$ and repeat this process sequentially until the target resolution $r_{e}$ is reached.
+Once $r_{e}$ is achieved, we output the resulting feature vectors, which can be used to calculate the FPD score.
+The parameters and modules corresponding to each resolution $r$ are listed in SI Sec.S1; Table S4. The feature extraction algorithm is outlined in Algorithm 1.
 Algorithm 1 Feature Extraction Algorithm
-1:  Input: Duration $r_{e}$, Start resolution $r_{s}$, Batch size $N$, Input data $x_{r_{s}}$, Feature dimension $D_{r_{s}}$, Data Length $L_{r_{s}}$ at resolution $r_{s}$
+1:Input: Duration $r_{e}$, Start resolution $r_{s}$, Batch size NN, Input data $x_{r_{s}}$, Feature dimension $D_{r_{s}}$, Data Length $L_{r_{s}}$ at resolution $r_{s}$
 2:Output: Final feature representation for data $x_{r_{s}}$ in $r_{e}$ scale
@@
 6:  else if $r_{s}$ is available then
-7:        Set $\mathbf{I}_{r_{s}}$ to $[0,x_{r_{s}}]$
+7:   Set $\mathbf{I}_{r_{s}}$ to $0,x_{r_{s}}]$
 8:  end if
@@
 Then, we can calculate the FPD scores based on the extracted feature vectors.
-For the architecture design of a single feature extractor in both models, we utilize residual blocks in the feature extractor as demonstrated in Fig.1.
-The residual connections have been widely studied and proved to increase neural network depth to enhance representation capabilities.
+For the architecture design of a single feature extractor in both models, we utilize residual blocks in the feature extractor as demonstrated in Fig.
+1.
+The residual connections have been widely studied and proven to increase neural network depth to enhance representation capabilities.
 We utilize the conv-1d and batchnorm-1d layers for accommodating the time-series inputs.
 This ResNet-like feature extractor is a base component for different scenarios and timescale tasks, followed by simple linear heads for different downstream training objectives.
-Figure 1: Overview of the hierarchical multi-resolution feature extraction framework, integrating multiple time-scale feature extractors to enhance generative model evaluation in power systems.
+Figure 1: Overview of the hierarchical multi-resolution feature extraction framework, integrating multiple time-scale feature extractors to enhance synthetic data evaluation in smart grids.
 II-C Fréchet Power-Scenario Distance (FPD) calculation
 Once robust representations are obtained from the feature extractor, we can then compare the distribution similarity between the two datasets in the feature space.
-Three distribution evaluation metrics have often been adopted[7]: (1) KL-divergence, (2) Kernel Distance (Maximum Mean Discrepancy - MMD), and (3) Fréchet Distance (Wasserstein-2 Distance).
-The features can often be approximately viewed as having a Gaussian distribution.
-Denote the feature vectors extracted from two sets $x_{r}^{1}$ and $x_{r}^{2}$ at resolution $r$ as $z_{r}^{1}$ and $z_{r}^{2}$.
-The means of extracted features can then be expressed as:
+Three distribution-level evaluation metrics are commonly used[20]: (1) Jensen–Shannon (JS) Divergence, (2) Kernel Distance (Maximum Mean Discrepancy - MMD), and (3) Fréchet Distance (Wasserstein-2 Distance).
+Denote the feature vectors extracted from two sets $x_{r}^{1}$ and $x_{r}^{2}$ at resolution $r$ as $z_{r}^{1}$ and $z_{r}^{2}$. The means of extracted features can then be expressed as:
 $\mathbf{m}_{1}=\frac{1}{N}\sum_{i=1}^{N}z_{r,i}^{1},\quad\mathbf{m}_{2}=\frac{1}{N}\sum_{i=1}^{N}z_{r,i}^{2}$ (2)
@@
 $\Sigma_{2}=\frac{1}{N}\sum_{i=1}^{N}\left(z_{r,i}^{2}-\mathbf{m}_{2}\right)\left(z_{r,i}^{2}-\mathbf{m}_{2}\right)^{T}$ (4)
-Using the KL-divergence as the evaluating metric, the similarity between $z_{r}^{1}$ and $z_{r}^{2}$ can be assessed as [15]:
-$\text{KL}(z_{r}^{1}||z_{r}^{2})=$ $\frac{1}{2}\bigg{(}\text{Tr}(\Sigma_{2}^{-1}\Sigma_{1})+(\mathbf{m}_{2}-\mathbf{m}_{1})^{T}\Sigma_{2}^{-1}(\mathbf{m}_{2}-\mathbf{m}_{1})$ (5)
-$-d+\log\left(\frac{\det(\Sigma_{2})}{\det(\Sigma_{1})}\right)\bigg{)},$
-where $d$ is the dimensionality of the feature space $z_{r}^{1}$ and $z_{r}^{2}$.
-It is important to note that when applying the KL-divergence to compare the distance between two normal distributions, both $z_{r}^{1}$ and $z_{r}^{2}$ must have the same dimensionality.
-Because our goal is to design a metric that can compare datasets in different resolutions which have different feature dimensions, the KL-divergence is not suitable for our purpose.
-It is also well known that the KL divergence has the disadvantage of being asymmetric and unstable in some cases.
+Using the Jensen–Shannon (JS) divergence—the symmetrized and smoothed variant of KL [21] —the similarity between $z_{r}^{1}$ and $z_{r}^{2}$ is defined as
+$\mathrm{JS}(z_{r}^{1}\,\|\,z_{r}^{2})$ $=\tfrac{1}{2}\,\mathrm{KL}\!\big(\mathcal{N}(\mathbf{m}_{1},\Sigma_{1})\,\big\|\,\mathcal{N}(\mathcal{M}\big)$ (5)
+$\quad+\tfrac{1}{2}\,\mathrm{KL}\!\big(\mathcal{N}(\mathbf{m}_{2},\Sigma_{2})\,\big\|\,\mathcal{N}(\mathcal{M}\big).$
+where $\mathcal{M}$ is the (Gaussian) mixture approximation with mean and covariance $\mathbf{m}_{M}=\tfrac{1}{2}(\mathbf{m}_{1}+\mathbf{m}_{2})$ and $\Sigma_{M}=\tfrac{1}{2}(\Sigma_{1}+\Sigma_{2})$, and $\mathrm{KL}(\cdot\|\cdot)$ denotes the closed-form KL divergence between Gaussians:
+$\mathrm{KL}\!\left(\mathcal{N}(\mathbf{m}_{1},\Sigma_{1})\,\middle\|\,\mathcal{N}(\mathbf{m}_{2},\Sigma_{2})\right)=\tfrac{1}{2}\Big[\operatorname{tr}(\Sigma_{2}^{-1}\Sigma_{1})+$ (6)
+$(\mathbf{m}_{2}-\mathbf{m}_{1})^{\top}\Sigma_{2}^{-1}(\mathbf{m}_{2}-\mathbf{m}_{1})-d+\log\!\frac{\det\Sigma_{2}}{\det\Sigma_{1}}\Big].$
+Here $d$ is the dimensionality of the feature space for $z_{r}^{1}$ and $z_{r}^{2}$.
+JS divergence is symmetric and bounded, offering improved stability over plain KL.
+However, when the distributions have little or no support overlap, JS becomes insensitive—it can saturate at its maximum and fail to track changes in similarity reliably, as illustrated in [22].
 With the Kernel distance as the evaluation metric, the distance between $z_{r}^{1}$ and $z_{r}^{2}$ is calculated as:
@@
 where $k(x,y)$ is a kernel function (e.g., Gaussian kernel).
-Choosing a proper kernel is extremely important for the metric to be effective, as different kernel functions can lead to varying metric performance.
-We propose to use the Fréchet Distance due to its stability and consistency as a metric based on our experiments.
-Since this metric is specifically tailored to evaluate power system scenarios, we name it Fréchet Power-Scenario Distance (FPD).
+Choosing a proper kernel is extremely important for the metric to be effective, as different kernel functions can lead to varying metric performances.
+We use the Fréchet distance due to its empirically observed stability and consistency.
+Since this metric is specifically tailored to evaluate power system scenarios, we refer to this as Fréchet Power-Scenario Distance (FPD).
 The $\text{FPD}(z_{r}^{1},z_{r}^{2})$ between the extracted features at resolution $r$ is calculated as:
 $\text{FPD}(z_{r}^{1},z_{r}^{2})=\|\mathbf{m}_{1}-\mathbf{m}_{2}\|^{2}+\text{Tr}(\Sigma_{1}+\Sigma_{2}-2(\Sigma_{1}\Sigma_{2})^{1/2})$ (8)
-FPD is essentially a form of the Fréchet Distance, where we assume that the features of the two datasets follow Gaussian distributions.
+FPD is essentially a form of the Fréchet distance, where we assume that the features of the two datasets follow Gaussian distributions [^1] For practical reasons, we consider only the mean and variance of distributions.
+Since the Gaussian maximizes entropy subject to fixed mean and covariance, we model feature distributions as multivariate Gaussian..
 This metric quantifies the similarity between the two distributions: a smaller FPD value indicates a more significant similarity between the feature distributions, ultimately reflecting a higher similarity between the datasets.
 III Model Training
+The purpose of training is to guide the feature extraction model to produce features that are tightly coupled to the smart grid datasets while retaining strong representational power.
 The training data used in this study consists of multiple open-source datasets, each capturing different generative tasks.
 These datasets include various types of demand, EV charging profiles, and various types of electricity generation.
-Each dataset varies in terms of resolution and timescale, contributing to a comprehensive representation of the power system domain.
-The specific characteristics of each dataset are summarized in Table II.
-TABLE II: Summary of Datasets Used for Training
-Dataset Description Source Resolution Timescale
-1 Wind Generation[16] NREL 5-minute 1 year
-2 Solar Generation[17] NREL 5-minute 1 year
-3 EV charging profiles[18] Northwest Power and Conservation Council Hourly 1 year
-4 Commercial user load[19] OpenEI.org Hourly 1 years
-5 Residential user load[19] OpenEI.org Hourly 1 years
-6 Load Data in Texas[20] ERCOT Hourly 3 years
-7 Demand by balancing authority[21] EIA daily 3 years
-8 Generation by balancing authority[21] EIA daily 3 years
-9 Utility Scale Electricity Net Generation[22] EIA monthly 14 years
-10 Phasor Measurement Unit PSSE simulations 0.016-second(60Hz) 50 hours
-For steady-state model, each module in the feature extraction model is trained individually and sequentially from the bottom level (5-min module) to the top level (yearly module).
-Once a certain level module has finished training, the parameters will be fixed, and we start to train the next level module.
-The training data for the module at each resolution is $I_{r_{s}}$ The training dataset is shown in Table II
+Each dataset varies in terms of resolution and timescale, contributing to a comprehensive representation of the smart grid domain.
+The specific characteristics of each dataset are summarized in SI Sec.S3; Table S8.
+For the steady-state model, each module in the feature extraction model is trained individually and sequentially from the bottom level (5-minute module) to the top level (yearly module).
+Once a certain level module has finished training, the parameters will be fixed, and we will start to train the next level module.
+The training data for the module at each resolution is $I_{r_{s}}$.
+The training dataset is shown in SI Sec.S3; Table S8.
 Each module is trained based on the loss function:
@@
 $f_{\text{clf}}(I_{r_{s}})=W_{\text{clf}}\mathcal{M}(I_{r_{s}})+b_{\text{clf}}$ (11)
-The first term in the (8), $\sum_{k=1}^{K}\left(f_{\text{reg}}^{k}(I_{r_{s}})-y_{i}^{\text{reg}}\right)^{2}$, corresponds to the Mean Squared Error (MSE) loss across a total of $K=9$ regression tasks.
-Each regression task is designed for leading extractors to capture features that reflect a specific statistical characteristic of the data, represented by $y_{i}^{\text{reg}}$.
-The list of regression tasks is shown in Table V in Appendix.
-The second term,$-\sum\limits_{i=1}^{C}y_{i}^{\text{clf}}\log\left(\text{softmax}(f_{\text{clf}}(I_{r_{s}}))\right)$, represents the Cross-Entropy Loss for the classification task which categorize data based on classification labels $y_{i}^{\text{clf}}$.
-$y_{i}^{\text{clf}}$ is determined by the dataset index shown in Table II and $C$ is the number of categories of dataset.
-The hyper-parameters for training are shown in Table VI in Appendix.
-For transient-state model training, we employ the fault type, amplitude minimum, and maximum as supervised pretraining labels.
-We then exploit the pre-trained feature extractor to evaluate synthetic samples.
-The length of a data sequence is 960, and input data is encoded into a 2048-dimension representation.
-Moreover, the hyper-parameters for training are shown in Table VII in Appendix.
-Once training is complete, the tasks are irrelevant, and the obtained module $\mathcal{M}(I_{r_{s}})$ will be used for feature extraction.
+[…]
+For transient-state model training, we use fault type, minimum amplitude, and maximum amplitude as supervised pretraining labels.
+We then exploit the pretrained feature extractor to evaluate synthetic samples.
+The length of a data sequence is 960, and the input data is encoded into a 2048-dimensional representation.
+Moreover, the hyper-parameters for training are shown in SI Sec.S1; Table S3.
+Once training is complete, we discard the task heads and use the trained module $\mathcal{M}(I_{r_{s}})$ solely for feature extraction.
 All training and experiments were conducted on a 16 GB Apple M2 chip MacBook Pro.
@@
 IV Effectiveness Evaluation of the FPD
-IV-A Effectiveness on different disturbances
-(a) Gaussian Noise
-(b) Missing Data
-(c) Solar Data Contaminate
-(d) Gaussian Smooth
-(e) Error Cumulate
-(f) Time Shift
-Figure 2: Effect of six different types of disturbances on the dataset $X$: (2(a)) Gaussian Noise is added with mean 0 and varying variances $\alpha=[0,0.16,1.6,4]$, where a larger $\alpha$ indicates more noise.
-(2(b)) Missing Data is simulated by randomly replacing $\alpha\times 100\%$ of the points in each sample of $X$ with zeros, where $\alpha=[0,0.1,0.25,0.5]$.
-(2(c)) Solar Data Contamination replaces $\alpha\times 100\%$ of the points in $X$ with corresponding points from a solar power dataset $Y$[17], where $\alpha=[0,0.25,0.5,0.75]$.
-(2(d)) Gaussian Smooth applies Gaussian filter which has variance $\alpha=[0,10,20,30]$ to the original dataset.
-Where a larger $\alpha$ indicates more significant smoothing.
-(2(e)) Error Cumulate is simulated by adding an error, which is times with a random Gaussian noise with a mean of 1 and varying variances $\alpha=[0,0.005,0.01,0.03]$ at each time-step.
-(2(f)) Time Shift is a potential problem that exists in time-series generation.
+IV-A Power-Aware Feature Extraction vs. Generic Encoders
+In this section, we demonstrate that the proposed extractor produces representations that align more closely with these physical principles than generic time-series encoders (e.g., TS2Vec [18], TimeMixer [19]) on solar generation data.
+For each extractor/encoder, we compute similarity using the same Fréchet distance in the respective feature space:
+  • Period Offset: Power signals exhibit strong periodicity (e.g., daily/seasonal load patterns; solar generation following the diurnal cycle—zero at night, peak near noon).
+To test sensitivity to disturbed periodic structure, we apply a time offset of the solar trace by $\alpha\in\{0,2,4\}$ hours.
+Here, $\alpha$ quantifies how far the shifted solar profile is displaced from its true period.
+A robust extractor/encoder should distinguish these shifted variants from the nominal series.
+  • Nighttime Generation (Physical Violation): Power signals must respect physical constraints (e.g., bounded ramp rates; no solar output at night).
+We inject nonzero solar generation during nighttime to create a controlled violation and evaluate whether extractors/encoders can detect it.
+The violation level is parameterized by $\alpha\in\{0,2,3\}$, where $\alpha$ denotes the number of nighttime hours forced to have nonzero solar output.
+  • Cross–Time-Resolution Consistency: A key property of power signals is that the same underlying process may be represented at different sampling resolutions; representations derived from hourly and 5-minute data should remain consistent.
+Our feature extractor employs a hierarchical, multi-scale design to capture structure across resolutions, whereas generic encoders (e.g., single-scale inputs) may treat different resolutions as unrelated sequences.
+To demonstrate this advantage, we compare features from hourly solar/wind datasets against their 5-minute counterparts and assess their similarity.
+As shown in Fig.
+2, the proposed extractor consistently and effectively distinguishes period shifts, detects physical feasibility violations, and preserves similarity judgments across time resolutions.
+By contrast, features from generic encoders do not reliably reveal phase misalignment or feasibility violations and yield inconsistent cross-resolution similarity, indicating reduced sensitivity to power-specific structure.
+These findings demonstrate that our extractor is better aligned with power-system characteristics than generic models.
+By faithfully capturing the distinct temporal regularities, physical constraints, and multi-resolution structure of power data, it provides a solid foundation for more reliable evaluation and analysis in subsequent sections.
+Figure 2: Features from the proposed power-aware extractor vs.
+generic time-series encoders, evaluated using the Fréchet distance in feature space in case: (a) Period Offset: solar series are shifted by $\alpha\in\{0,2,4\}$ hours.
+(b) Solar Physical Violation: Unrealistic power generation is injected for $\alpha\in\{0,2,3\}$ hours of solar generation during night.
+(c) Cross Time Resolution: 5-minute solar compared against hourly solar/wind.
+IV-B Distributional Metric Suitability in Learned Feature Space
+Figure 3: Effectiveness of FPD with six different types of disturbances on the dataset $X$: (a) Gaussian Noise is added with mean 0 and varying variances $\alpha=[0,0.16,1.6,4]$, where a larger $\alpha$ indicates more noise.
+(b) Missing Data is simulated by randomly replacing $\alpha\times 100\%$ of the points in each sample of $X$ with zeros, where $\alpha=[0,0.1,0.25,0.5]$.
+(c) Solar Data Contamination replaces $\alpha\times 100\%$ of the points in $X$ with corresponding points from a solar power dataset YY[23], where $\alpha=[0,0.25,0.5,0.75]$.
+(d) Gaussian Smooth applies Gaussian filter which has variance $\alpha=[0,10,20,30]$ to the original dataset.
+A larger $\alpha$ indicates more significant smoothing.
+(e) Error Cumulate is simulated by adding an error, which is multipled by a random Gaussian noise with a mean of 1 and varying variances $\alpha=[0,0.005,0.01,0.03]$ at each time step.
+(f) Time Shift is a potential problem that exists in time-series generation.
 To simulate this effect, we shift the data forward by $\alpha=[0,40,60,80]$ intervals.
-In this section, we conduct several experiments to demonstrate the effectiveness of FPD.
-The experiments are designed based on the Wind Generation dataset 1 in Table II.
-We progressively increase the disturbance applied to the original dataset X in each experiment to provide a controllable and scalable measure of FPD performance.
-  • Gaussian Noise: In this case, Gaussian noise with a mean of 0 and a variance $\alpha$ is added to $X$.
-Where $\alpha\in[0,0.16,1.6,4]$ represents the disturbance added.
-Larger $\alpha$ indicates more noise added.
+In the previous section, we showed that the learned feature space produced by our extractor aligns with key smart grid properties.
+Here, we further present experiments that motivate the choice of the Fréchet distance over alternative distributional metrics for measuring similarity within this learned space.
+The experiments are designed based on the Wind Generation Dataset (Dataset 1 in SI Sec.S3; Table S8) We progressively increase the disturbance applied to the original dataset X in each experiment to provide a controllable and scalable measure of FPD performance.
+  • Gaussian Noise: In this case, Gaussian noise with a mean of 0 and a variance $\alpha$ is added to $X$, where $\alpha\in[0,0.16,1.6,4]$ denotes the disturbance level.
+A larger $\alpha$ indicates more noise added.
   • Missing Data: To simulate missing data points in $X$, we randomly replace some values with zeros for each sample in this case.
 The disturbance $\alpha\in[0,0.1,0.25,0.5]$ indicates that $\alpha\times 100\%$ of the points in $X$ are randomly set to zero.
-  • Solar Data Contamination: To show FPD can distinguish the data from another dataset, we contaminate the wind dataset $X$ with data from the solar power dataset $Y$ [17].
+  • Solar Data Contamination: To test cross-dataset discrimination, we contaminate the wind dataset $X$ with samples from the solar dataset $Y$ [23].
 The disturbance $\alpha\in[0,0.25,0.5,0.75]$ indicates that $\alpha\times 100\%$ of the samples in $X$ are replaced by corresponding points from $Y$.
@@
   • Time Shift: Time shift in time-series data is a common issue in generative models, often leading to misalignment of temporal features between the original and generated data.
-This misalignment can be particularly problematic in power system data, where certain temporal features, such as peaks, carry critical information.
+This misalignment can be particularly problematic in smart grid data, where certain temporal features, such as peaks, carry critical information.
 To simulate this, we shift the data forward by intervals of $\alpha=[0,40,60,80]$, creating misleading temporal dependencies.
@@
 As shown in Fig.
-2, KL-divergence can be inconsistent due to its asymmetry feature, while Kernel Distance may fail to capture significant changes and is sensitive to kernel choice.
+3, JS-divergence is neither consistent nor effective across different disturbance events, while Kernel distance may fail to capture significant changes and is sensitive to kernel choice.
 Additionally, both methods also exhibit scale variability across disturbances.
@@
 For these reasons, we select Fréchet distance as the evaluation metric and refer to our method as Fréchet Power-Scenario Distance (FPD).
+IV-C Comparison of FPD with conventional evaluation metrics
 Figure 4: Sample-wise distance MAPE vs. FPD
-IV-B Comparison of FPD with conventional evaluation metrics
-Figure 4: Data-level Fréchet distance (FD) VS. FPD
-IV-B1 Sample-Wise Euclidean Distance (MAPE) vs. Distribution-Wise Distance (FPD)
-Traditional Euclidean-based metrics such as Mean Absolute Percentage Error (MAPE) typically perform sample-wise comparisons, making the results highly dependent on the selection of sample pairs between the two datasets.
+Figure 5: Data-level distribution metrics vs. FPD
+IV-C1 Sample-Wise Euclidean Distance (MAPE) vs. Distribution-Wise Distance (FPD)
+Traditional Euclidean-based metrics, such as Mean Absolute Percentage Error (MAPE), typically perform sample-wise comparisons, making the results highly dependent on the selection of sample pairs between the two datasets.
 This approach can lead to inconsistent and potentially invalid outcomes.
 To illustrate this, we revisit the Gaussian noise case presented in Fig.
-2(a).
-In this example, we make MAPE to match sample pairs randomly between the original and noise-perturbed datasets.
+3.
+In this example, we compute MAPE after randomly matching sample pairs between the original and noise-perturbed datasets.
 As shown in Fig.
-3, in some instances, higher noise levels result in lower MAPE values, which contradicts the ground truth.
+4, in some instances, higher noise levels result in lower MAPE values, which contradicts the ground truth.
 In contrast, the proposed FPD metric demonstrates expected behavior, providing a more effective measurement.
-IV-B2 Data-Level Distribution Comparison (Fréchet Distance) vs. Feature-Level Distribution Comparison (FPD)
-As previously noted, the data-level Fréchet Distance (FD) metric is often ineffective at capturing high-level complex patterns.
-FPD metric addresses this issue by incorporating feature extraction and measuring the similarity between distributions in the feature space.
-Feature extraction enables a better representation of higher-level patterns and structures within the data, allowing FPD to capture more relevant similarities for evaluating the quality of generated data.
-To demonstrate this advantage, we constructed a fabricated dataset by randomly sampling from a distribution with a similar mean and covariance to the original data.
-Despite these similarities, the fabricated dataset lacks the temporal patterns present in the original dataset, resulting in a sparse relation.
+IV-C2 Data-Level Distribution Comparison vs. Feature-Level Distribution Comparison (FPD)
+Although distribution-level scoring rules such as CRPS and the Energy Score, as well as distributional metrics like JS-divergence, MMD, and Wasserstein distance, avoid the sample-wise limitations discussed above, they still struggle to detect changes in generation quality when applied directly to raw data.
+In particular, such data-level metrics often underemphasize higher-order structure—such as temporal dependencies and cross-feature interactions.
+The proposed FPD metric addresses this gap by introducing a feature-extraction stage prior to distributional comparison: data are first mapped into a learned representation, and similarity is then evaluated in this feature space.
+This feature-based approach highlights higher-level patterns and structures, yielding similarity estimates that more faithfully capture the quality of generated data.
+To illustrate, we construct a fabricated dataset by randomly sampling from a distribution with mean and covariance matched to the original data.
+Despite this alignment in first- and second-order statistics, the fabricated dataset lacks the temporal structure of the original, resulting in weaker dependencies.
 As shown in Fig.
-4, while FD fails to identify this disparity, FPD accurately reflects the weaker relationship between the fabricated and original datasets.
-IV-C Coherence between FPD and the downstream tasks
-To assess whether FPD’s indication of high-quality data can be used to determine the practical usefulness for downstream tasks, we further evaluate its efficacy for transient-state case using PMU data containing three types of faults.
-We employ the Brownian noises to mimic the accumulating temporary errors in the PMU data prediction as depicted in Fig.
-5.
-As shown in Table III, we cross-evaluate different fault scenarios.
-Hence, we can observe non-zero diagonal elements in the FPD score matrix, along with the diagonal values being significantly lower than the non-diagonal elements, showing the capability of identifying different scenarios other than evaluating generation quality.
-To further prove the capability of FPD can handle the downstream task in steady-state case, we reuse the wind dataset[16] and define the sub-task as data classification based on the average power across the day as used in Chen’s work [23].
-The three cases differ based on average power levels, as shown in Fig.
-6.
-Moreover, the cross-evaluation of different mean value scenarios is shown in the Table.IV again highlights the alignment between the FPD evaluation and the specific sub-task.
-Figure 5: Ground-truth and synthetic PMU voltage magnitude curves under different fault scenarios. Synthetic data is modulated using Brownian noises.
-TABLE III: Cross-evaluation on FPD performance for PMU data.
-Synthetic $|V|$ Ground-truth $|V|$ bus_trip branch_trip branch_fault
-bus_trip 744.26 1086.07 2586.15
-branch_trip 1462.24 846.03 2720.04
-branch_fault 2516.35 2401.13 687.99
-Figure 6: Ground-truth and synthetic wind power under different mean scenarios. Synthetic data is modulated using Brownian noises.
-TABLE IV: Cross-evaluation on FPD performance for wind power data.
-Synthetic $|MW|$ Ground-truth $|MW|$ $\bar{x}<0.5$ $\bar{x}<3$ $\bar{x}>6$
-$\bar{x}<0.5$ 736.97 1300.42 901.36
-$\bar{x}<3$ 1300.42 565.35 2720.04
-$\bar{x}>6$ 901.36 989.83 686.51
-IV-D Cross-model comparison with FPD
+5, CRPS and the Energy Score show little separation between the fabricated dataset and a lightly perturbed Gaussian baseline, while the Fréchet distance also fails to flag the discrepancy.
+By contrast, FPD correctly identifies a weaker relationship between the fabricated and original datasets.
+IV-D Robustness of FPD to Labeling Schemes and Sample Budgets for Downstream Tasks
+We ask whether FPD’s indication of data quality reliably predicts downstream utility under realistic variation in labeling and sampling.
+In practice, given a downstream task, generative models can synthesize effectively unbounded scenario sets whose usefulness depends on both the scenario definition and the sampling strategy.
+We therefore test whether FPD remains aligned with downstream performance as labeling schemes change and the sampling budget varies.
+Specifically, we evaluate FPD under alternative wind-power ramp labeling schemes and across multiple sample budgets.
+We use the SDWPF dataset [24] with 10-minute resolution and reshape each example as a 1-hour window of six 10-minute values per turbine ($X\in\mathbb{R}^{N\times 6}$). Let the ramp rate be
+$r=\frac{P_{50}-P_{00}}{P_{\max}},$
+where $P_{00}$ and $P_{50}$ are the first and last 10-minute values within the hour.
+We label SDWPF ramp events by direction and magnitude using two threshold sets inspired by [25] and [26] as shown in Table I.
+TABLE I: Ramp-rate category definitions for the two labeling scenarios.
+Category Scenario 1 (coarser) Scenario 2 (finer)
+Strong down-ramp r<0.50r<-0.50 r<0.30r<-0.30
+Moderate down-ramp $0.50\leq r<-0.33$ $0.30\leq r<-0.20$
+Mild down-ramp $0.33\leq r<-0.25$ $0.20\leq r<-0.10$
+Neutral $0.25\leq r\leq 0.25$ $0.10\leq r\leq 0.10$
+Mild up-ramp $0.25<r\leq 0.33$ $0.10<r\leq 0.20$
+Moderate up-ramp $0.33<r\leq 0.50$ $0.20<r\leq 0.30$
+Strong up-ramp $r>0.50$ $r>0.30$
+We then train two Wasserstein GAN (W-GAN) generators: W-GAN-full on the complete dataset (all categories) and W-GAN-collapse after removing all up-ramping categories to induce mode collapse (the model learns only down-ramp patterns).
+For each scenario, an identical MLP classifier is trained on real data labeled by that scenario and applied to synthetic samples from each generator at three sizes (0.2k/2k/20k).
+We report weighted-F1 and macro-F1.
+Weighted-F1 is the weighted average F1 score, so it reflects performance on prevalent categories and is a proxy for the perceived quality of typical generations.
+Macro-F1 is the unweighted mean of the F1 score, so it treats all categories equally and is sensitive to diversity.
+As shown in Table II, both generators achieve similarly high weighted F1, which indicates high-quality generation on prevalent categories, whereas macro-F1 drops markedly for W-GAN-collapse, revealing it is ignoring up-ramp categories, which is a classic mode collapse.
+To compute the FPD score, we first upsample the real and synthetic series to 5-minute resolution, extract hourly features, and then measure the Fréchet distance between the generated and real feature distributions.
+Across both scenarios and all sample sizes, W-GAN-full consistently attains lower (better) FPD than W-GAN-collapse, with values stable as sample size increases.
+These results demonstrate that FPD aligns with the downstream classification diagnosis and remains robust to the choice of labeling scenario as well as sampling size.
+TABLE II: Classification and FPD (scaled by $10^{-4}$) results under two scenarios
+Scenario 1 Scenario 2
+W-GAN-full W-GAN-collapse W-GAN-full W-GAN-collapse
+Weighted F1 Macro F1 FPD Weighted F1 Macro F1 FPD Weighted F1 Macro F1 FPD Weighted F1 Macro F1 FPD
+Sample_0.2k 0.994 0.955 454.4 0.987 0.640 903.5 0.995 0.936 454.4 0.989 0.679 794.3
+Sample_2k 0.998 0.984 549.1 0.980 0.602 923.3 0.994 0.918 549.1 0.989 0.673 731.7
+Sample_20k 0.997 0.980 550.7 0.984 0.631 885.8 0.994 0.946 550.7 0.990 0.676 800.5
+IV-E Cross-model comparison with FPD
 In the previous sections, all evaluations were conducted directly on the artificially generated datasets.
 This approach enables precise control over generation quality and model performance assessment.
-However, FPD can also be used to evaluate model generated performing the same task, as demonstrated in the following two scenarios.
+However, FPD can also be used to evaluate model generation performance on the same task, as demonstrated in the following two scenarios.
   1. 1.
-    Scenario 1: EV Charging Profile Data Generation
-    Accurately evaluating the quality of synthetic data in real-world settings presents a significant challenge, as highlighted by [24], for which FPD provides a solution.
-[24] proposed a novel generative model, DiffCharge, based on a diffusion model designed to conditionally generate electric vehicle (EV) charging power profiles.
-Probability density function (PDF) comparisons and discriminative scores derived from a trained classifier were used to evaluate battery-level data generation performance between DiffCharge and baseline models such as Time-GAN.
-However, evaluations for station-level data generation were not given.
-We conduct experiments to compare DiffCharge’s performance with a W-GAN model on the station level; see Fig.
-7.
-We find that DiffCharge achieves a lower FPD score.
-Visually, we also observe that DiffCharge can capture the general trend better, while W-GAN generates less realistic peaky signals.
-Therefore, our results indeed support the conclusion that the DiffCharge model can more accurately approximate the distribution of EV charging profiles at both the station-level and battery-level.
+    Case 1: EV Charging Profile Data Generation
+    In [27], the authors present DiffCharge, a diffusion-based generative model that combines an LSTM backbone with multi-head self-attention to conditionally generate EV charging power profiles.
+They show that DiffCharge better captures short-/long-range temporal dependencies and cross-time interactions than GAN baselines for both battery-level (single-EV sessions) and station-level (aggregate daily load) scenarios.
+Building on this line of work—and because station-level evaluations were not reported in the underlying study—we train a Wasserstein GAN (W-GAN) baseline on station-level data, compare it against DiffCharge, and assess both using our FPD metric.
+We also report Kernel distance (RBF) and CRPS.
+As shown in Fig.
+6, DiffCharge better preserves general temporal structure, whereas W-GAN produces unrealistic, long-lasting peaks.
+Consistently, FPD yields larger distances for W-GAN outputs—aligning with Kernel distance and CRPS—while offering even more apparent discrimination.
+Hyperparameter choices for DiffCharge and W-GAN appear in SI Sec.S2; Table S5 and Table S7, respectively.
   2. 2.
-    Scenario 2: Wind Power Data Generation
-    To further demonstrate FPD’s cross-model comparison ability, we replicate the study from [1], where the authors introduced a model called StyleGAN, which integrates forecast and climate data to enhance the generation of renewable energy data.
-They showed that StyleGAN is superior to WGAN proposed in [23] based on multiple uncertainty quantification and statistical metrics, such as RMSE.
-In our experiment, the differences between generations from StyleGAN and WGAN were challenging to distinguish through visual inspection.
-StyleGAN’s results seemed to capture the overall trend more accurately.
+    Case 2: Wind Power Data Generation
+    To further demonstrate FPD’s cross-model comparison ability, we replicate the study of [1], which proposed a StyleGAN that integrates conditions on meteorological variables (e.g., temperature, humidity, weather codes) and employs a sequence encoder to infer day-ahead forecast patterns, yielding more realistic generation than GAN baselines such as the conditional W-GAN of [28].
+In our replication, we randomly sample day-ahead meteorology paired with the corresponding day-ahead forecast as conditioning inputs and generate 1-day wind-power trajectories (96 15-minute intervals) under a mixed weather scenario.
+As a baseline, we train a conditional W-GAN with the same conditioning and sampling protocol to generate the same horizon.
 As shown in Fig.
-8, FPD confirms StyleGAN aligns better with real data, supporting the findings in [1].
-    Figure 7: Scenario 1 – Comparison of DiffCharge and W-GAN in generating EV charging profiles at the station level.
-    Figure 8: Scenario 2 – Comparison of W-GAN and StyleGAN in generating wind power data.
-IV-E Comparability across varying time resolutions
-A key feature of FPD is its ability to adapt to varying time resolutions within a single metric.
-This flexibility enables effective comparisons across datasets with different time resolutions.
-To demonstrate this capability, we collected hourly resolution data for solar and wind power and evaluated their similarity with the 5-minute resolution solar and wind power datasets used in Section IV-A.
-The results, shown in Fig.
-9, indicate that FPD successfully identifies that the solar power datasets are more similar to each other than they are to the wind power data despite their differing time resolutions.
-This result demonstrates the metric’s effectiveness in facilitating comparability across datasets with different temporal resolutions.
-Figure 9: The FPD metric is evaluated between the left:5-min resolution solar power dataset[17], right top:hourly resolution solar power dataset[25] and right bottom:hourly resolution wind power dataset[25].
-The results demonstrate that FPD effectively generalizes across different time resolutions.
+7, samples from StyleGAN track the overall trend and variability of the real data more closely, whereas W-GAN—despite being meteorology-aware—fails to capture the realistic structure in this case.
+This ranking is correctly reflected by our proposed FPD metric and by CRPS, while a kernel distance computed directly on the raw data fails to distinguish the models.
+These results underscore the importance of the feature extraction design in FPD.
+Hyperparameter choices for StyleGAN and W-GAN appear in SI Sec.S2; Table S6 and Table S7, respectively.
+    Figure 6: Case 1 – Comparison of DiffCharge and W-GAN in generating EV charging profiles at the station level.
+    Figure 7: Case 2 – Comparison of W-GAN and StyleGAN in generating wind power data.
 V Conclusion and Future Work
-This paper introduces the Fréchet Power-Scenario Distance (FPD) as a novel, task-agnostic metric for evaluating generative model outputs in the power systems sector.
-FPD addresses the limitations of traditional metrics by employing a hierarchical feature extraction framework that captures multi-resolution temporal information.
-Then, the Fréchet Distance is applied to assess the quality of generated data in the feature space.
-This method enables FPD to provide an accurate, robust, and consistent evaluation across diverse power system applications.
-Comprehensive case studies demonstrate FPD’s effectiveness in distinguishing data quality across various disturbances, time resolutions, and sub-task scenarios.
-These results underscore FPD’s potential as a standardized cross-model approach for validating generative models, offering support for both enhancing model quality and promoting wider adoption within the power systems domain.
-There are some potential studies for future work.
+This paper introduces the Fréchet Power-Scenario Distance (FPD)—a feature-based, distributional metric tailored to smart grid data for evaluating synthetic datasets.
+FPD encodes both generic temporal structure and domain-specific properties before computing distances in the learned feature space.
+This design yields consistent, effective judgments across disturbances, time resolutions, and scenario definitions, and better aligns with downstream utility, where raw-data metrics and generic encoders often fall short.
+Comprehensive studies demonstrate FPD’s effectiveness in distinguishing data quality across various disturbances, time resolutions, and sub-task scenarios.
+These results underscore FPD’s potential as a standardized cross-model approach for validating synthetic datasets, offering support for both enhancing model quality and promoting wider adoption within the smart grid domain.
+There are several directions for future work.
 1) Currently, the dataset we are using is still limited due to accessibility constraints.
-Thus, building a more comprehensive dataset tailored to a wider range of tasks within the power system domain is possible to improve the performance of FPD; 2) While the current model structure can handle multi-resolution and multi-duration data, it is not yet sufficiently generalized to accommodate data of any resolution or duration.
+Thus, building a more comprehensive dataset tailored to a broader range of tasks within the smart grid domain could improve the performance of FPD; 2) While the current model structure can handle multi-resolution and multi-duration data, it is not yet sufficiently general to accommodate data of any resolution or duration.
 Therefore, there is potential for developing a more generalized structure that can address this challenge and further enhance the versatility of the metric.
-Appendix A Appendix: Feature extraction training detail
-TABLE V: Regression labels for steady-state training
-Regression Tasks
-Minimum
-Maximum
-Range
-Variance
-Number of data points above average
-Number of data points below average
-Skewness
-Kurtosis
-Coefficient of variation
-TABLE VI: Hyper-parameters for steady-state model training
-Parameter Setting
-Minibatch size 1024
-Optimizer Adam
-Optimizer: learning rate 1e-3
-TABLE VII: Hyper-parameters for transient-state model training
-Parameter Setting
-Minibatch size 128
-Optimizer Adam
-Optimizer: learning rate 1e-4
-Optimizer: weight decay 2e-5
-Classification loss coefficient 1.0
-Regression loss coefficient 1.0